Seventy-nine percent of adults in a certain country believe that life on other planets is plausible. You randomly select five adults and ask them whether they believe that life on other planets is plausible. The random variable represents the number of adults who believe that life on other planets is plausible. Find the​ mean, variance, and standard deviation of the binomial distribution for the random variable. Interpret the results.

Find the mean of the binomial distribution.
Answer #1

The random variable ( X ) represents the number of adults who believe that life on other planets is plausible.

79 % of adults in a certain country believe that life on other planets is plausible.

p = 79% = 0.79

five adults are randomly selected , n = 5

X follows Binomial distribution with n = 5 and p = 0.79

If X follows Binomial distribution with parameter n and p then ,

Mean = np

Variance = np(1-p)

Standard deviation = Vnp(1-0)

Mean = np= 5*0.79 = 3.95

On​ average, 3.95 out of every 5 adults in the country believe that life on other planets is plausible.

Variance = np(1-p) = 5*0.79*(1-0.79) = 0.8295

Standard deviation = Vnp(1-0) = 5*0.79(1 -0.79)    = 0.9108